OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2.

OpenStudy (anonymous):

I think f(5)=17 but im not sure

OpenStudy (anonymous):

@freckles sorry

OpenStudy (freckles):

f(1)+f(2)=2+3=5 right?

OpenStudy (freckles):

so techically you have f(n)=5+f(n-1) correct?

OpenStudy (freckles):

all this is saying to find the nth term you will need to add 5 to the previous term

OpenStudy (anonymous):

yeah, f(3)=5 f(4)=10 and f(5)=12 right?

OpenStudy (anonymous):

no f(5)= 15

OpenStudy (freckles):

well the second term is 3 and 3+5 will be the third term

OpenStudy (freckles):

all you are doing is adding 5 to the previous term to find the term

OpenStudy (anonymous):

oh so 2,3,8,13,18

OpenStudy (freckles):

sounds good in slow motion: 

OpenStudy (anonymous):

and 13+8=18

OpenStudy (freckles):

:)

OpenStudy (freckles):

you got this